# Localize-and-Stitch: Efficient Model Merging via Sparse Task Arithmetic

### v1

@@ Appendix C Details on localization
$\operatornamewithlimits{argmin}_{\gamma\in\{0,1\}^{{d}}:\|\gamma\|_{o}\leq s}\ell_{\tau}(\gamma\odot\theta_{\text{ft}}+(1-\gamma)\odot\theta_{\text{pre}})$
They also use a reparametriztion of the binary mask $\gamma$ as the sigmoid of a real-valued vector $S$, and reformulate the problem as
$\operatornamewithlimits{argmin}_{S\in\mathbb{R}^{d}}\ell_{i}(\gamma\odot\theta_{\text{ft}}+(1-\gamma)\odot\theta_{\text{pre}}),$
@@
$S_{i}=\operatornamewithlimits{argmin}_{S\in\mathbb{R}^{d}}\ell_{i}\left(\theta_{\text{pre}}+\sigma(S)\odot\tau_{i}\right)+\lambda\|\sigma(S)\|_{1},$
There are two main differences between the formulations. Firstly, our formulation of $S$ is more straightforward, as we directly have $\gamma=\sigma(S)$. In contrast, $S$ in Appendix C serves as a selector of whether to take the value from $\gamma_{base}$, leading to more complex computation. Secondly, our approach uses the $L_{1}$ constraint to control the sparsity in a more fine-grained manner, while Appendix C does not have this constraint, and the authors control the sparsity via early stopping.
Table 13: Comparison of localization methods on RoBERTa-base models on twelve language tasks.
Task SST-2 CR MR MPQA TREC SUBJ QNLI SNLI MNLI RTE MRPC QQP Average
@@
Single-task grafted [45] 0.902 0.908 0.862 0.851 0.884 0.925 0.752 0.756 0.676 0.643 0.757 0.693 0.801
@@
Task SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Average
@@
Single-task grafted [45] 0.731 0.750 0.935 0.959 0.929 0.932 0.976 0.753 0.871
@@
Merged [45] 0.669 0.678 0.798 0.861 0.846 0.826 0.919 0.653 0.781
We present the performance comparison of the two localization methods in Tables 13 and 14. In both cases, our approach with Equation 1 outperforms [45]. The performance may come from the fact that [45] use early stopping to control the sparsity, which results in incomplete optimization for the masks. We also report the merged performance by following the same stitching process. On the language tasks, the performance is similar, while on the vision tasks, our localization leads to better merged performance.
Appendix D GPT2-XL experiment details
@@ Appendix E Datasets
Following the practice in [30], we use the following 8 datasets for the vision part of our experiments:
  • SUN397 [66]. The Scene UNderstanding dataset contains 108,754 images of 397 classes.
  • Stanford Cars [33]. The Stanford Cars dataset contains 16,185 images of 196 classes of cars.
  • RESISC45 [6]. The REmote Sensing Image Scene Classification dataset contains 31,500 images, covering 45 scene classes.
  • EuroSAT [25]. The EuroSAT dataset consists of 10 classes with 27000 labeled and geo-referenced samples. Each class represents a different land use and land cover.
  • SVHN [41]. The Street View House Numbers dataset contains 600,000 digit images in 10 classes of printed digits cropped from pictures of house number plates.
  • GTRSB [53]. The German Traffic Sign Recognition Benchmark contains 43 classes of traffic signs with more than 50,000 images.
  • MNIST [34]. The MNIST dataset contains 60,000 training images and 10,000 testing images of 10 handwritten digits.
  • DTD [8]. The Describable Texture Dataset contains 5,640 texture images in the wild with 47 classes.
@@
Following the practice in [45], we use the following 12 datasets for the language part of our experiments. The majority comes from the GLUE benchmark [59].
  • SST-2 [52]. The Stanford Sentiment Treebank is a sentiment analysis dataset, which contains sentences from movie reviews and human annotated binary sentiments.
  • CR [28]. The Customer Review dataset consists of customer reviews on e-commerce products with binary sentiment labels.
  • MR [44]. The Movie Review dataset consists of movie reviews with binary sentiment labels.
  • MPQA [61]. The Multi-Perspective Question Answering dataset contains news articles and text documents manually annotated for opinions and other private states including beliefs, emotions, sentiments, etc. Here, we use it for binary sentiment classification.
  • TREC [58]. The Text REtrieval Conference (TREC) dataset contains 6k questions phrased by users and categorized into a small number of categories. The task is to classify the questions into these categories.
  • SUBJ [43]. The SUBJectivity dataset contains 10k movie reviews with an annotation of whether the review describes something subjective or objective about the movie.
  • QNLI [59]. The Question-answering NLI dataset is converted from the Stanford Question Answering Dataset (SQuAD) [51], which contains questions and the paragraphs that contain the answer to the corresponding questions. QNLI converts SQuAD into sentence pair classification by forming a pair between each question and each sentence in the corresponding context, where the task is to predict whether the context contains the answer to the question.
  • SNLI [2]. The Stanford Natural Language Inference dataset contains 570k sentence pairs manually labeled as entailment, contradiction or neutral.
  • MNLI [62]. The Multi-Genre Natural Language Inference Corpus is a collection of 433k sentence pairs annotated with textual entailment information.
  • RTE [59]. The Recognizing Textual Entialment dataset contains a series of textual entailment challenges, including RTE1 [11], RTE2 [21], RTE3 [18] and RTE5 [1]. The neutral and contradiction classes are combined into a no entailment class.
  • MRPC [15]. The Microsoft Research Paraphrase Corpus consists of sentence pairs from online news sources, with human annotations of whether the sentences in the pair are semantically equivalent. Since the classes are imbalanced, we report the F1 score.
  • QQP [31]. The Quora Question Pairs dataset consists of question-answer pairs from the website Quora. The task is to determine whether two questions are semantically equivalent.
@@
  • MMLU [26]. The Massive Multitask Language Understanding measures knowledge in 57 subjects across STEM, humanities, social science, etc.
  • ARC [9]. The AI2 Reasoning Challenge dataset contains 7,787 grade-school level, multiple choice science questions.
  • TruthfulQA [36]. The TruthfulQA dataset measure whether a model is truthful in generating answers to questions. It comprises 817 questions spanning 38 categories, including health, law, finance, etc.
  • Alpaca [56]. The Alpaca dataset contains 52,000 instruction-following examples.
  • GSM8k [10]. The Grade School Math 8K dataset contains 8,500 high quality grade school math problems created by human problem writers. The problems take 2 to 8 stpes to solve.
  • HotpotQA [69]. The HotpotQA dataset is a question answering dataset featuring multi-hop questions.
@@
Appendix F Implementation details
@@
Finetuning. For the experiments on RoBERTa-base, we perform the finetuning process following the same procedure as [45]. Specifically, we use a batch size of 4 and a learning rate of 2e-5 to finetune on each of the language tasks for 10 epochs with the SGD optimizer. For the experiments on CLIP ViT, we directly use the finetuned checkpoints provided in [30] with the data preprocessing step provided in [68]. The finetuned models in the GPT2-XL experiments in provided in Appendix D.
Localization. Following the practice in [45], in the localization step, we initialize the trainable real-valued vector $S$ as the mask for top-$k\%$ largest entries in the task vector. Since the actual map is rounded from $\sigma(S)$ but not $S$, we choose the initial values of $S$ to be either 0 or 3, as $\sigma(3)$ is sufficiently close to 1. To achieve a sparsity level of $1\%$, we use the learning rate $71e$ 7, batch size 16, $L_{1}$ regularization factor $\lambda$ 1e-5 and perform the optimization for 10 epochs on 64-shot data from each task. Following common practice in [45, 67], we only perform localization in the transformer blocks, and do not consider embedding layers.
@@
Appendix G Limitations
Our work shares some limitations with existing model merging methods. Firstly, the applicability of our method is limited to merging models finetuned from the same initialization. Secondly, although in Section 3.1, we provide empirical findings that support our motivations, it remains unclear whether the arguments can be tested theoretically. In fact, the theoretical understanding of the effectiveness of model merging is still underexplored, only with the few attempts from linear mode connectivity [17] and tangent task spaces [42]. Thirdly, although the conflicts among tasks can be significantly reduced with our approach, the performance still lags behind the single-task finetuned models. This phenomenon calls for more understanding of the difficulties in model merging beyond task conflicts.

### v2

@@ Appendix E Details on localization
$\operatorname*{arg\,min}_{\gamma\in\{0,1\}^{{d}}:\|\gamma\|_{o}\leq s}\ell_{\tau}(\gamma\odot\theta_{\text{ft}}+(1-\gamma)\odot\theta_{\text{pre}})$
They also use a reparametrization of the binary mask $\gamma$ as the sigmoid of a real-valued vector $S$, and reformulate the problem as
$\operatorname*{arg\,min}_{S\in\mathbb{R}^{d}}\ell_{i}(\gamma\odot\theta_{\text{ft}}+(1-\gamma)\odot\theta_{\text{pre}}),$
@@
$S_{i}=\operatorname*{arg\,min}_{S\in\mathbb{R}^{d}}\ell_{i}\left(\theta_{\text{pre}}+\sigma(S)\odot\tau_{i}\right)+\lambda\|\sigma(S)\|_{1},$
There are two main differences between the formulations. Firstly, our formulation of $S$ is more straightforward, as we directly have $\gamma=\sigma(S)$. In contrast, $S$ in Appendix E serves as a selector of whether to take the value from $\gamma_{base}$, leading to more complex computation. Secondly, our approach uses the $L_{1}$ constraint to control the sparsity in a more fine-grained manner, while Appendix E does not have this constraint, and the authors control the sparsity via early stopping.
Table 16: Accuracy (F1) comparison of localization methods on RoBERTa-base models on twelve language tasks.
Task SST-2 CR MR MPQA TREC SUBJ QNLI SNLI MNLI RTE MRPC QQP Average acc/F1
@@
Single-task grafted (Panigrahi et al., 2023) 0.902 0.908 0.862 0.851 0.884 0.925 0.752 0.756 0.676 0.643 0.757 0.693 0.801
@@
Task SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Average acc
@@
Single-task grafted Panigrahi et al. (2023) 0.731 0.750 0.935 0.959 0.929 0.932 0.976 0.753 0.871
@@
Merged Panigrahi et al. (2023) 0.669 0.678 0.798 0.861 0.846 0.826 0.919 0.653 0.781
We present the performance comparison of the two localization methods in Tables 16 and 17. In both cases, our approach with Equation 1 outperforms Panigrahi et al. (2023). The performance may come from the fact that Panigrahi et al. (2023) use early stopping to control the sparsity, which results in incomplete optimization for the masks. We also report the merged performance by following the same stitching process. On the language tasks, the performance is similar, while on the vision tasks, our localization leads to better merged performance.
Appendix F GPT2-XL experiment details
@@ Appendix G Datasets
Following the practice in Ilharco et al. (2023), we use the following 8 datasets for the vision part of our experiments:
  • SUN397 (Xiao et al., 2016). The Scene UNderstanding dataset contains 108,754 images of 397 classes.
  • Stanford Cars Krause et al. (2013). The Stanford Cars dataset contains 16,185 images of 196 classes of cars.
  • RESISC45 (Cheng et al., 2017). The REmote Sensing Image Scene Classification dataset contains 31,500 images, covering 45 scene classes.
  • EuroSAT (Helber et al., 2019). The EuroSAT dataset consists of 10 classes with 27000 labeled and geo-referenced samples. Each class represents a different land use and land cover.
  • SVHN (Netzer et al., 2011). The Street View House Numbers dataset contains 600,000 digit images in 10 classes of printed digits cropped from pictures of house number plates.
  • GTRSB (Stallkamp et al., 2011). The German Traffic Sign Recognition Benchmark contains 43 classes of traffic signs with more than 50,000 images.
  • MNIST (LeCun et al., 2010). The MNIST dataset contains 60,000 training images and 10,000 testing images of 10 handwritten digits.
  • DTD (Cimpoi et al., 2014). The Describable Texture Dataset contains 5,640 texture images in the wild with 47 classes.
@@
Following the practice in Panigrahi et al. (2023), we use the following 12 datasets for the language part of our experiments. The majority comes from the GLUE benchmark (Wang et al., 2018).
  • SST-2 (Socher et al., 2013). The Stanford Sentiment Treebank is a sentiment analysis dataset, which contains sentences from movie reviews and human annotated binary sentiments.
  • CR (Hu & Liu, 2004). The Customer Review dataset consists of customer reviews on e-commerce products with binary sentiment labels.
  • MR (Pang & Lee, 2005). The Movie Review dataset consists of movie reviews with binary sentiment labels.
  • MPQA (Wiebe et al., 2005). The Multi-Perspective Question Answering dataset contains news articles and text documents manually annotated for opinions and other private states including beliefs, emotions, sentiments, etc. Here, we use it for binary sentiment classification.
  • TREC (Voorhees et al., 1999). The Text REtrieval Conference (TREC) dataset contains 6k questions phrased by users and categorized into a small number of categories. The task is to classify the questions into these categories.
  • SUBJ (Pang & Lee, 2004). The SUBJectivity dataset contains 10k movie reviews with an annotation of whether the review describes something subjective or objective about the movie.
  • QNLI (Wang et al., 2018). The Question-answering NLI dataset is converted from the Stanford Question Answering Dataset (SQuAD) (Rajpurkar et al., 2016), which contains questions and the paragraphs that contain the answer to the corresponding questions. QNLI converts SQuAD into sentence pair classification by forming a pair between each question and each sentence in the corresponding context, where the task is to predict whether the context contains the answer to the question.
  • SNLI (Bowman et al., 2015). The Stanford Natural Language Inference dataset contains 570k sentence pairs manually labeled as entailment, contradiction or neutral.
  • MNLI (Williams et al., 2017). The Multi-Genre Natural Language Inference Corpus is a collection of 433k sentence pairs annotated with textual entailment information.
  • RTE (Wang et al., 2018). The Recognizing Textual Entialment dataset contains a series of textual entailment challenges, including RTE1 (Dagan et al., 2005), RTE2 (Haim et al., 2006), RTE3 (Giampiccolo et al., 2007) and RTE5 (Bentivogli et al., 2009). The neutral and contradiction classes are combined into a no entailment class.
  • MRPC Dolan & Brockett (2005). The Microsoft Research Paraphrase Corpus consists of sentence pairs from online news sources, with human annotations of whether the sentences in the pair are semantically equivalent. Since the classes are imbalanced, we report the F1 score.
  • QQP (Iyer et al., ). The Quora Question Pairs dataset consists of question-answer pairs from the website Quora. The task is to determine whether two questions are semantically equivalent.
@@
  • MMLU (Hendrycks et al., 2021). The Massive Multitask Language Understanding measures knowledge in 57 subjects across STEM, humanities, social science, etc.
  • ARC (Clark et al., 2018). The AI2 Reasoning Challenge dataset contains 7,787 grade-school level, multiple choice science questions.
  • TruthfulQA (Lin et al., 2021). The TruthfulQA dataset measure whether a model is truthful in generating answers to questions. It comprises 817 questions spanning 38 categories, including health, law, finance, etc.
  • Alpaca (Taori et al., 2023). The Alpaca dataset contains 52,000 instruction-following examples.
  • GSM8k (Cobbe et al., 2021). The Grade School Math 8K dataset contains 8,500 high quality grade school math problems created by human problem writers. The problems take 2 to 8 stpes to solve.
  • HotpotQA (Yang et al., 2018). The HotpotQA dataset is a question answering dataset featuring multi-hop questions.
@@
Appendix H Implementation details
@@
Finetuning. For the experiments on RoBERTa-base, we perform the finetuning process following the same procedure as Panigrahi et al. (2023). Specifically, we use a batch size of 4 and a learning rate of 2e-5 to finetune on each of the language tasks for 10 epochs with the SGD optimizer. For the experiments on CLIP ViT, we directly use the finetuned checkpoints provided in Ilharco et al. (2023) with the data preprocessing step provided in (Yang et al., 2023). The finetuned models in the GPT2-XL experiments in provided in Appendix F.
Localization. Following the practice in Panigrahi et al. (2023), in the localization step, we initialize the trainable real-valued vector $S$ as the mask for top-$k\%$ largest entries in the task vector. Since the actual map is rounded from $\sigma(S)$ but not $S$, we choose the initial values of $S$ to be either 0 or 3, as $\sigma(3)$ is sufficiently close to 1. To achieve a sparsity level of $1\%$, we use the learning rate $71e$ 7, batch size 16, $L_{1}$ regularization factor $\lambda$ 1e-5 and perform the optimization for 10 epochs on 64-shot data from each task. Following common practice in Panigrahi et al. (2023); Yadav et al. (2023), we only perform localization in the transformer blocks, and do not consider embedding layers.